OpenStudy (mathhelp346):

When Gauss was very young, a teacher told him to find the sum of all the numbers from 1-100, thinking this would keep him busy. Gauss cleverly paired the numbers and observed that 1+99=100, 2+98=100, 3+97=100, and so on.

OpenStudy (mathhelp346):

How many pairs of numbers can be formed this way? shoe me how you got it too pleaseeeeeeeeeeee

OpenStudy (amistre64):

50

OpenStudy (mathhelp346):

really? how

OpenStudy (amistre64):

100/2 = 50 pairs

OpenStudy (amistre64):

err.. lol 100 pairs ... im getting aheadf of myself

OpenStudy (mathhelp346):

are you sure? there's an answer sheet and it says 49 pairs..

OpenStudy (mathhelp346):

do you know they got it?

OpenStudy (amistre64):

lets start small if need be: 1,2,3,4,5,6 ; how many pairs can you make similar to the above question?

OpenStudy (amistre64):

6,1 5,2 4,3 3,4 2,5 1,6 6 pairs total

OpenStudy (amistre64):

but you may notice that you have counted twice as many times as need be when you go that route; so we have to divide it by half

OpenStudy (mathhelp346):

what i did was 99,1 98,2 97,3 ......

OpenStudy (amistre64):

guass observed that the first and last numbers added; not the first and next to last numbers

OpenStudy (mathhelp346):

..?

OpenStudy (mathhelp346):

isn't what you just said the same??

OpenStudy (amistre64):

it still works with an odd count to begin with; but only cause the math is forgiving :)

OpenStudy (amistre64):

100+1 = 101 99+2 = 101 98+3 = 101 ......

OpenStudy (mathhelp346):

??

OpenStudy (amistre64):

101, 100 times = 10100; but youve counted each set twiice, so divide in half to get 5050

OpenStudy (mathhelp346):

how'd you get that

OpenStudy (amistre64):

the same way i got it with the smaller example

OpenStudy (amistre64):

n*(first+last)/2

OpenStudy (mathhelp346):

n=??

OpenStudy (amistre64):

i spose if you start counting with 99+1, itll be 49; but that just seems wierd

OpenStudy (mathhelp346):

what does n equal?

OpenStudy (amistre64):

n = number of terms in the set

OpenStudy (amistre64):

1,2,3,...,98,99,100 ; has 100 terms in the set right?

OpenStudy (mathhelp346):

yea

OpenStudy (mathhelp346):

so...100*(first+last)/2 ??

OpenStudy (amistre64):

so you count 100 pairs of (101); 100+1 = 101 99+2 = 101 98+3 = 101 -1+1 = 101

OpenStudy (amistre64):

100(100+1) = 10100 now tell me, does 1+100; and 100+1 count as seperate or the same?

OpenStudy (mathhelp346):

OH i understand that, but how did the answer book get 49 for the pairs of numbers?

OpenStudy (mathhelp346):

it's the same, right

OpenStudy (mathhelp346):

??

OpenStudy (amistre64):

the started counting off at 99+1 instead of doing it sensibly

OpenStudy (mathhelp346):

so how would i show work for that? would i have to write down the numbers 1-100??

OpenStudy (amistre64):

the sum of all the numbers from 1-99 is 99(99+1)/2; which is not the same

OpenStudy (amistre64):

if your teacher wants you to show all the work instead of just an understanding of it, then yes; youd write them all out

OpenStudy (mathhelp346):

...aw

OpenStudy (amistre64):

the thing is you shouldnt have to start at 100 to understand the concept ...

OpenStudy (amistre64):

its the same concept for 2 numbers as it is for 100 or 100000

OpenStudy (amistre64):

it even works for 1 number :)

OpenStudy (mathhelp346):

whaaa..?

OpenStudy (amistre64):

what is the sum of 1?

OpenStudy (mathhelp346):

umm 1?

OpenStudy (amistre64):

yep; so if we count it twice; what do we get? 1+1 = ?

OpenStudy (mathhelp346):

2..

OpenStudy (mathhelp346):

what does this all mean??

OpenStudy (amistre64):

and how do we get it back to its proper sum? divide by 2 right?

OpenStudy (amistre64):

lets try it with 2 numbers, 1,2 ; what is the sum of the numbers?

OpenStudy (mathhelp346):

3

OpenStudy (amistre64):

now what do we get when we sum up pairs? 2+1 = 3 1+2 = 3 -------- 6 right?

OpenStudy (mathhelp346):

yea

OpenStudy (amistre64):

but you can see that we simply counted twice as much as we needed to right?

OpenStudy (mathhelp346):

yea

OpenStudy (mathhelp346):

but you're not suppose to count twice

OpenStudy (mathhelp346):

wait is that why you divided by 2?

OpenStudy (amistre64):

Yes :)

OpenStudy (mathhelp346):

ok

OpenStudy (amistre64):

by counting pairs, we actually end up counting it all 2 times; so we have to adjust for that

OpenStudy (mathhelp346):

so how would i show my work? the sum of the numbers 1-100 is 5050

OpenStudy (mathhelp346):

i don't want to take up too much time and space, so what would be the easiest way?

OpenStudy (amistre64):

you would either have to show that all the pairs of numbers from 1 to 100 is double the actual sum; or, devise a plan from recursive equations

OpenStudy (mathhelp346):

..??

OpenStudy (amistre64):

without some basic background in recurrsives; the long way is about the only way i can see to prove it

OpenStudy (mathhelp346):

oh

OpenStudy (amistre64):

it either that or prove it for the smaller cases and then generalize it for larger cases

OpenStudy (mathhelp346):

how would i show that all the pairs of numbers from 1-100 is double the actual sum though

OpenStudy (mathhelp346):

how did you get 10100 anyways?

OpenStudy (amistre64):

Show that 1+100 = 100+1 id say.

OpenStudy (amistre64):

101*100 = 10100

OpenStudy (amistre64):

i counted up the 100 pairs that summed to 101

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

you want to know how it is made 100*(101)/2

OpenStudy (mathhelp346):

uhh??

OpenStudy (amistre64):

1 + 2 + 3 + ... + 98 + 99 + 100 100 + 99 + 98 + ... + 3 + 2 + 1 ---------------------------------- 101+101+101 + ... +101+101+ 101

OpenStudy (mathhelp346):

ok i understand now

OpenStudy (mathhelp346):

if i don't understand can i call u in the chat..?

OpenStudy (amistre64):

i dont go to chat alot, jamal might tho

OpenStudy (mathhelp346):

who..? nvm thanks :D

OpenStudy (mathhelp346):

OHH jamal..

OpenStudy (anonymous):

I want to tell you another method that is very powerfull by which you can know sum of squares, cubes what so ever

OpenStudy (mathhelp346):

ok

OpenStudy (anonymous):

x^2-(x-1)^2=2x-1 1^2-(0)^2=2(1)-1 2^2_1^2=2(2)-1 ..................... (n-1)^2-(n-2)^2=2(n-1)-1 n^2-(n-1)^2=2n-1 adding all the terms you will see lot of terms are delting n^2=2(1+2+3....n)-(1+1+1+1.....n) n^2=2(1+2+3....n)-n so 1+2+3...n=n(n+1)/2